### v4

@@ 2.2.3 Theoretical guarantees
With the proper parameter choices, KRILL is guaranteed to solve any restricted KRR problem with any positive-semidefinite kernel matrix $\bm{A}\in{}^{N\times N}$ and any regularization parameter $\mu>0$. Unlike RPCholesky precondioning, KRILL does not require any eigenvalue decay. Here, we state our main theoretical guarantee, which is proved in Section 5.2.
@@
Fix a failure probability $\delta\in(0,1)$ and an error tolerance $\varepsilon\in(0,1)$. Let $\bm{A}$ be any positive-semidefinite matrix; let $\mathsf{S}$ be an index set of cardinality $|\mathsf{S}|=k$; and define $\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$. Draw a random sparse sign embedding $\bm{\Phi}\in{}^{d\times N}$ with column sparsity $\zeta$ and embedding dimension $d$ satisfying
$\zeta=\mathcal{O}(\log(k/\delta))\quad\text{and}\quad d=\mathcal{O}(k\log(k/\delta)).$ (2.10)
With probability at least $1-\delta$ over the randomness in $\bm{\Phi}$, the KRILL preconditioner $\bm{P}=\bm{A}(\mathsf{S},:)\bm{\Phi}^{\ast}\bm{\Phi}\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ controls the condition number at a level
@@
Conditional on the event (2.11), when we apply preconditioned CG to the linear system $\bm{M}\widehat{\bm{\beta}}=\bm{A}(\mathsf{S},:)\bm{y}$, we obtain an approximate solution $\widehat{\bm{\beta}}^{(t)}$ for which
@@
at any iteration $t\geq\log(2/\varepsilon)$, where $\widehat{\bm{\beta}}$ is the actual solution.
Theorem 2.3 implies that KRILL can solve any restricted KRR problem in $\mathcal{O}((N+k^{2})k\log k)$ operations. In contrast, FALKON-type preconditioners [44, 30, 45] are only guaranteed to solve restricted KRR problems if the number of centers and regularization satisfy $k=\Omega(\sqrt{N})$ and $\mu=\Omega(\sqrt{N})$ [44, 45]. Given these constraints, it is no surprise that KRILL performs more robustly than FALKON in our experiments.
Figure 6 evaluates KRILL’s capacity to control the conditioning of the 20 example problems when $\mu/N=10^{-12}$. Before KRILL preconditioning, the condition numbers range from $10^{1}$$10^{15}$. After KRILL preconditioning, the condition numbers are concentrated in the range $10^{1}$$10^{2}$. The preconditioned eigenvalues approximately align with one another, and they align with the inverse square singular values of a $2k\times k$ Gaussian matrix, which has the approximate condition number $(1+(1/\sqrt{2})^{2}/(1-(1/\sqrt{2}))^{2}\approx 34.0$ (for intuition, see [37, Sec. 2]). This comparison shows that the performance of KRILL is largely independent of the details of the kernel matrix $\bm{A}$ and only depends on the embedding dimension $d=2k$.
Figure 6: Eigenvalue decay for the 20 KRR problems in Table 1. Left panel shows normalized eigenvalues of the matrix $\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ before KRILL preconditioning (normalized by the largest eigenvalue). Right panel shows normalized eigenvalues of the matrix $\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}$ after KRILL preconditioning.
@@ 3 Background and comparisons with other preconditioners
In this section, we compare our new preconditioners with existing preconditioners for solving the full-data and restricted KRR equations. To begin, we observe that the full-data and restricted KRR equations both take the form
@@
where $\bm{M}\in{}^{d\times d}$ is a positive-definite matrix and $\bm{b}\in{}^{d}$ is a vector. In full-data KRR, the dimension $d$ equals the number of data points, $d=N$. In restricted KRR, the dimension equals the number of data centers, $d=k$.
Preconditioned conjugate gradient
The conjugate gradient (CG) algorithm [46, §6.7] is a popular approach for solving linear systems of the form (3.1). When $\bm{M}$ is well-conditioned ($\kappa(\bm{M})\approx 1$), CG gives a high-accuracy solution using only a small number of iterations, with each iteration requiring a matrix–vector product with $\bm{M}$. More precisely, the following bound [46, eq. (6.128)] controls the convergence rate of the CG iterates $\bm{z}^{(0)},\bm{z}^{(1)},\ldots$ to the solution $\bm{z}$ of the system (3.1):
@@
The full-data and restricted KRR equations are typically ill-conditioned, so we need to apply a positive-definite preconditioner $\bm{P}\in{}^{d\times d}$ to these problems to improve the convergence. The preconditioned CG algorithm (Algorithm 3, described in [24, §10.3]), is equivalent to applying standard CG to the preconditioned system
@@
1:Subroutine Product computing the action $\bm{u}\mapsto\bm{M}\bm{u}$, subroutine Preconditioner computing the action $\bm{u}\mapsto\bm{P}^{-1}\bm{u}$, right-hand-side vector $\bm{b}\in{}^{N}$, and tolerance $\varepsilon$.
@@
In the rest of this section, we review the history of preconditioned CG methods for the full-data and restricted KRR equations. We do not cover KRR methods based on sketch-and-solve [6, 1] or stochastic gradient descent [17, 28], since the accuracy of these methods is more limited [29, §10].
@@ 3.1 Preconditioners for full-data KRR
where $\bm{\widehat{A}}\in{}^{N\times N}$ is an approximation of the full kernel matrix $\bm{A}$. Ideally, we would obtain an accurate approximation $\bm{\widehat{A}}\approx\bm{A}$, and we would quickly apply $\bm{P}^{-1}$ to a vector at every CG step.
In the KRR literature, $\bm{\widehat{A}}$ is typically constructed as a low-rank approximation of $\bm{A}$ [16, 5, 23, 53, 22]. A low-rank approximation can only be accurate when the kernel matrix is close to being low-rank, i.e., when the eigenvalues of $\bm{A}$ decay quickly. However, even under favorable eigenvalue decay conditions, it has been a challenge to identify accurate and computationally tractable low-rank approximations.
A popular low-rank approximation method for KRR problems is the column Nyström approximation [29, §19.2], which is obtained from a partial Cholesky decomposition with pivoting. The column Nyström approximation takes the form
@@
where $\textsf{S}=\{s_{1},s_{2},\ldots,s_{r}\}$ coincides with the pivots chosen in the Cholesky procedure, and it lists the columns of $\bm{A}$ that participate in the approximation. The column Nyström approximation is convenient because it is formed from the columns of $\bm{A}$ indexed by S and does not require viewing the rest of the matrix. Yet the accuracy of the column Nyström approximation depends on selecting a good index set $\mathsf{S}$. In the context of full-data KRR, there are two prominent strategies for picking the set $\mathsf{S}$ of columns in the Nyström approximation, called uniform sampling and greedy selection. Both of these approaches exhibit serious failure modes.
Failure of uniform sampling
The first strategy selects the column pivots $s_{1},\ldots,s_{r}$ uniformly at random [16, 22]. This uniform sampling method can be effective for some problems, but it fails to explore less populated regions of data space. To illustrate this shortcoming, let $\bm{1}_{m}$ denote the $m$-vector with entries equal to $1$, and consider the kernel matrix
@@
In principle, constructing a preconditioner for $\bm{A}$ should be easy, since the $\mu$-tail rank (Definition 2.1) is just $\textup{rank}_{\mu}(\bm{A})=2$ for any $\mu<N^{1/3}$. However, uniform sampling selects many columns from the left block and neglects columns from the right block that are also needed to ensure the approximation quality. For uniform sampling to build an effective preconditioner, we need an approximation rank much higher than $2$ and at least as large as $r=\Omega(N^{2/3})$.
Failure of greedy selection
A second strategy for choosing columns is the greedy selection method [23, 53], in which we adaptively select each column pivot $s_{i}$ by finding the largest diagonal element of the residual matrix $\bm{A}-\bm{\widehat{A}}_{(i-1)}$ after $i-1$ steps of the Cholesky procedure. The greedy method has the opposite failure mode from the uniform method: it focuses on outlier data points and fails to explore highly populated regions of data space. To see this limitation, let $\mathbf{I}_{N^{2/3}}$ denote the $N^{2/3}\times N^{2/3}$ identity matrix, and consider the kernel matrix
[…]
for a small parameter $\delta>0$. Constructing a preconditioner for $\bm{A}$ is easy in theory, since $\textup{rank}_{\mu}(\bm{A})\leq 2$ for any $\mu\geq N^{2/3}\delta$. Yet the greedy strategy selects columns from the right block first and misses the left columns. To provide an effective preconditioner, greedy sampling needs an approximation rank $r>N^{2/3}$, which is large enough that all the columns in the right block have been selected.
@@
1:Positive semidefinite matrix $\bm{A}\in{}^{N\times N}$, approximation rank $k$, block size $B$.
2:Factor matrix $\bm{F}\in{}^{N\times k}$, index set $\mathsf{S}$.
@@
8:     $\bm{R}\leftarrow\textsc{Cholesky}(\bm{G}(\mathsf{S}^{\prime},:))$triangleright $\bm{G}(\mathsf{S}^{\prime},:)=\bm{R}^{*}\bm{R}$
@@
Advantages of RPCholesky
In this paper, we propose a new KRR preconditioner that uses RPCholesky (see Algorithm 4 or [12]) to select the columns for the Nyström approximation. In RPCholesky, we adaptively sample the column pivot $s_{i}$ with probability proportional to the diagonal entries of the residual matrix $\bm{A}-\bm{\widehat{A}}_{(i-1)}$ after $i-1$ steps of the Cholesky procedure. Because of the adaptive sampling distribution, RPCholesky balances exploration of the small diagonal entries and exploitation of the large diagonal entries, avoiding the failure modes of both the uniform and greedy strategies. RPCholesky is guaranteed to produce an effective preconditioner if the approximation rank $r$ is set to a modest multiple of the $\mu$-tail rank (Theorem 2.2).
Ridge leverage score sampling
Ridge leverage score sampling [1, 36, 45] is a more elaborate procedure for selecting the pivots in the Nyström approximation. This method first approximates the ridge leverage scores, which are the diagonal entries $\bm{\ell}^{\lambda}=\operatorname{diag}(\bm{A}(\bm{A}+\lambda\mathbf{I})^{-1})$ for a parameter $\lambda>0$. Then, the method randomly generates a Nyström approximation whose pivots are selected with probabilities proportional to the ridge leverage scores. While RLS sampling can lead to an approximation that is nearly as accurate as RPCholesky (see Figure 1), it can also produce lower-quality approximations given challenging inputs [12]. Additionally, RLS sampling is slower than the blocked implementation of RPCholesky given in Algorithm 4, because of the requirement to approximate the ridge leverage scores. For example, when we apply RLS sampling (with the RecursiveRLS algorithm [36]) to compute a rank-$1000$ approximation of a kernel matrix of size $N=4\times 10^{4}$, the runtime is 2.3×2.3$\times$2.3 × slower than RPCholesky’s runtime.
More expensive Gaussian preconditioner
@@
$\bm{\widehat{A}}=\bm{A}\bm{X}(\bm{X}^{*}\bm{A}\bm{X})^{\dagger}\bm{X}^{*}\bm{A},$
where $\bm{X}\in{}^{N\times r}$ is a matrix with independent standard normal entries. Forming the Gaussian Nyström approximation requires $r$ matrix–vector multiplications with the full kernel matrix, which is much more expensive than forming a column Nyström approximation (3.3) in factored form. To maintain a cost of $\mathcal{O}(N^{2})$ operations, we can only run Gaussian Nyström with a constant approximation rank $r=\mathcal{O}(1)$, whereas we can run RPCholesky preconditioning with a larger approximation rank $r=\mathcal{O}(\sqrt{N})$. Because of this larger approximation rank, RPCholesky preconditioning leads to a more accurate approximation with stronger theoretical guarantees.
Other approaches
We briefly mention two other types of approximations that can be used to build preconditioners for the full-data KRR problem. First, certain types of kernel matrices can be approximated using random features [16, 5], but the quality of the approximation tends to be poor. Numerical tests indicate that the random features approach does not yield a competitive preconditioner [16, 55]. See Figures 1 and 2 for further numerical results. Second, the kernel matrix can also be approximated using a hierarchical low-rank approximation [16, 3, 11, 56]. The hierarchical approach has mainly been applied to low-dimensional input data, and the applicability to high-dimensional data sets remains unclear [3].

### v5

@@ 2.2.3 Theoretical guarantees
With the proper parameter choices, KRILL is guaranteed to solve any nonsingular restricted KRR problem. Unlike RPCholesky precondioning, KRILL does not require any eigenvalue decay. Here, we state our main theoretical guarantee, which is proved in Section˜5.2.
@@
Fix a failure probability $\delta\in(0,1)$ and an error tolerance $\varepsilon\in(0,1)$. Let $\bm{A}$ be any positive-semidefinite matrix, let $\mathsf{S}$ be an index set of cardinality $|\mathsf{S}|=k$, and let $\mu>0$ be a positive number. Define $\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ and assume $\bm{M}$ is nonsingular. Generate a sparse random sign embedding $\bm{\Phi}\in\real^{d\times N}$ where the column sparsity $\zeta$ and embedding dimension $d$ are functions of $k$ and $\delta$ given in [16] that satisfy
$\zeta=\mathcal{O}(\log(k/\penalty 50\delta))\quad\text{and}\quad d=\mathcal{O}(k\log(k/\penalty 50\delta)).$ (2.10)
With probability at least $1-\delta$, the KRILL preconditioner $\bm{P}=\bm{A}(\mathsf{S},:)\bm{\Phi}^{\ast}\bm{\Phi}\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ controls the condition number at a level
@@
Conditional on the event (2.11), when we apply preconditioned CG to the linear system $\bm{M}\widehat{\bm{\beta}}=\bm{A}(\mathsf{S},:)\bm{y}$, we obtain an approximation $\widehat{\bm{\beta}}^{(t)}$ to the actual solution $\widehat{\bm{\beta}}$ that satisfies
@@
at any iteration $t\geq\log(2/\penalty 50\varepsilon)$.
Theorem˜2.3 implies that KRILL can solve any nonsingular restricted KRR problem in $\mathcal{O}((N+k^{2})\,k\log k)$ operations. In contrast, FALKON and related preconditioners [47, 34, 48] are only guaranteed to solve restricted KRR problems when the number of centers and the regularization are functions of $k$ satisfying $k=\Omega(\sqrt{N})$ and $\mu=\Omega(\sqrt{N})$ [47, 48]. Given these constraints, it is no surprise that KRILL performs more robustly than FALKON in our experiments.
Figure˜6 evaluates KRILL’s impact on the 20 example problems with a tiny regularization $\mu=10^{-12}N$. Before KRILL preconditioning, the condition number $\lambda_{\max}/\lambda_{\min}$ ranges from $10^{8}$$10^{15}$ (left panel). After KRILL preconditioning, the preconditioned eigenvalues (middle panel) are closely modeled by the inverse square singular values of a $2k\times k$ Gaussian matrix (right panel), leading to a condition number $(1+1/\sqrt{2})^{2}/(1-1/\sqrt{2})^{2}\approx 34.0$. This Gaussian comparison was previously presented in [40, Sec. 2]; and it is an example of the Gaussian universality phenomenon in random matrix theory [54, 14]. The comparison suggests that KRILL performs consistently for all kernel matrices $\bm{A}$, with performance only depending on the embedding dimension $d=2k$.
No preconditioner       KRILL eigenvalues     Eigenvalues $\lambda_{i}(\bm{Z}^{*}\bm{Z})^{-1}$
eigenvalues $\lambda_{i}(\bm{M})$       $\lambda_{i}(\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2})$      for $\bm{Z}\sim\mathcal{N}(0,1)^{2k\times k}$
Figure 6: Eigenvalue decay for the 20 KRR problems in Table˜1. Left panel shows eigenvalues of the matrix $\bm{M}=\bm{A}(\mathsf{S},:)\bm{A}(:,\mathsf{S})+\mu\bm{A}(\mathsf{S},\mathsf{S})$ before KRILL preconditioning (normalized by the largest eigenvalue). Middle panel shows normalized eigenvalues of the matrix $\bm{P}^{-1/2}\bm{M}\bm{P}^{-1/2}$ after KRILL preconditioning. Right panel shows eigenvalues $\lambda_{i}(\bm{Z}^{*}\bm{Z})^{-1}$ for a Gaussian matrix $\bm{Z}\sim\mathcal{N}(0,1)^{2k\times k}$. Note the different vertical axis scales.
@@ 3 Background and comparisons with other preconditioners
In this section, we compare our new conjugate gradient (CG) preconditioners with existing preconditioners for solving the full-data and restricted KRR equations. To begin, we observe that the full-data and restricted KRR equations both take the form
@@
where $\bm{M}\in\real^{d\times d}$ is a strictly positive definite matrix and $\bm{b}\in\real^{d}$ is a vector. In full-data KRR, the dimension $d$ equals the number of data points, $d=N$. In restricted KRR, the dimension equals the number of data centers, $d=k$.
Preconditioned conjugate gradient.
The CG algorithm [49, §6.7] is a popular approach for solving linear systems of the form (3.1). When $\bm{M}$ is well-conditioned ($\kappa(\bm{M})\approx 1$), CG gives a high-accuracy solution in just a small number of iterations, with each iteration requiring a matrix–vector product with $\bm{M}$. More precisely, the following bound [49, eq. 6.128] controls the convergence rate of the CG iterates $\bm{z}^{(0)},\bm{z}^{(1)},\ldots$ to the solution $\bm{z}$ of the system (3.1):
@@
The full-data and restricted KRR equations are typically ill-conditioned, so we need to apply a positive-definite preconditioner $\bm{P}\in\real^{d\times d}$ to these problems to improve the convergence. The preconditioned CG algorithm (Algorithm˜3, described in [27, §10.3]), is equivalent to applying standard CG to the preconditioned system
@@
1:Subroutine Product computing the action $\bm{u}\mapsto\bm{M}\bm{u}$, subroutine Preconditioner computing the action $\bm{u}\mapsto\bm{P}^{-1}\bm{u}$, right-hand-side vector $\bm{b}\in\real^{N}$, and tolerance $\varepsilon$.
@@
In the rest of this section, we review the history of preconditioned CG methods for the full-data and restricted KRR problems. The section does not cover KRR methods based on sketch-and-solve [6, 1] or row access methods like stochastic gradient descent [18, 32], since the paper’s focus is obtaining the best preconditioned CG method.
@@ 3.1 Preconditioners for full-data KRR
where $\bm{\widehat{A}}\in\real^{N\times N}$ is an approximation of the full kernel matrix $\bm{A}$. Ideally, the approximation $\bm{\widehat{A}}\approx\bm{A}$ is highly accurate, and we can quickly apply $\bm{P}^{-1}$ to a vector.
In the KRR literature, $\bm{\widehat{A}}$ is typically constructed as a low-rank approximation [17, 5, 26, 57, 25]. A low-rank approximation can only be accurate when the kernel matrix is close to being low-rank, i.e., when the eigenvalues of $\bm{A}$ decay quickly. However, even under favorable eigenvalue decay conditions, it has been a challenge to identify accurate and computationally tractable low-rank approximations.
One low-rank approximation used in KRR problems is the column Nyström approximation [33, §19.2], which is obtained from a partial Cholesky decomposition with pivoting. The column Nyström approximation takes the form
@@
where $\textsf{S}=\{s_{1},s_{2},\ldots,s_{r}\}$ is the set of pivots chosen in the Cholesky procedure. The column Nyström approximation is convenient because it is formed from the columns of $\bm{A}$ indexed by S and does not require viewing the rest of the matrix. Yet the approximation accuracy depends on the index set $\mathsf{S}$. In the context of full-data KRR, the most popular strategies for picking the set $\mathsf{S}$ of columns are uniform sampling and greedy selection. Both of these approaches exhibit failure modes.
Failure of uniform sampling.
One simple strategy is to select the column pivots $s_{1},\ldots,s_{r}$ uniformly at random [17, 25]. This uniform sampling method can be effective for some problems, but it fails to explore less populated regions of data space. To illustrate this shortcoming, let $\bm{1}_{m}$ denote the $m$-vector with entries equal to $1$, and consider the kernel matrix
@@
Constructing a low-rank approximation for $\bm{A}$ should be easy, since the rank is $2$. However, uniform sampling selects many columns from the left block and neglects columns from the right block that are also needed to ensure the approximation quality. Uniform sampling needs an approximation rank much higher than $2$ and at least as large as $r=\Omega(N^{2/3})$ to build an effective preconditioner.
Failure of greedy selection.
A second strategy for choosing columns is the greedy selection method [26, 57], in which we adaptively select each column pivot $s_{i}$ by finding the largest diagonal element of the residual matrix $\bm{A}-\bm{\widehat{A}}_{(i-1)}$ after i1i-1 steps of the Cholesky procedure. The greedy method has the opposite failure mode from the uniform method: it focuses on outlier data points and fails to explore highly populated regions of data space. To see this limitation, let $\mathbf{I}_{N^{2/3}}$ denote the $N^{2/3}\times N^{2/3}$ identity matrix, and consider the kernel matrix
$\bm{A}=\bm{1}_{N}^{\vphantom{\ast}}\bm{1}_{N}^{\ast}+\begin{pmatrix}\frac{1}{2N}\,\bm{1}_{N-N^{2/3}}^{\vphantom{*}}\bm{1}_{N-N^{2/3}}^{\ast}&~\\ &\frac{1}{N}\,\mathbf{I}_{N^{2/3}}\end{pmatrix}$.
Constructing a preconditioner for $\bm{A}$ should be easy, since the $\mu$-tail rank (Definition˜2.1) is $\textup{rank}_{\mu}(\bm{A})\leq 2$ for any $\mu\geq N^{-1/3}$. Yet the greedy strategy selects columns from the right block first and misses the left columns. To provide an effective preconditioner, greedy sampling needs an approximation rank $r>N^{2/3}$, which is large enough that all the columns in the right block have been selected.
@@
1:Positive semidefinite matrix $\bm{A}\in\real^{N\times N}$, approximation rank $k$, block size $B$.
2:Factor matrix $\bm{F}\in\real^{N\times k}$, index set $\mathsf{S}$.
@@
8:  $\bm{R}\leftarrow\textsc{Cholesky}(\bm{G}(\mathsf{S}^{\prime},:))$ $\triangleright$ $\bm{G}(\mathsf{S}^{\prime},:)=\bm{R}^{*}\bm{R}$
@@
Advantages of RPCholesky.
In this paper, we propose a new KRR preconditioner that uses RPCholesky (see Algorithm˜4 or [13]) to select the columns for the Nyström approximation. In RPCholesky, we adaptively sample the column pivot $s_{i}$ with probability proportional to the diagonal entries of the residual matrix $\bm{A}-\bm{\widehat{A}}_{(i-1)}$ after i1i-1 steps of the Cholesky procedure. Because of the adaptive sampling distribution, RPCholesky balances exploration of the small diagonal entries and exploitation of the large diagonal entries, avoiding the failure modes of both the uniform and greedy strategies. RPCholesky is guaranteed to produce an effective preconditioner if the approximation rank $r$ is set to a modest multiple of the $\mu$-tail rank (Theorem˜2.2).
Ridge leverage score sampling.
Ridge leverage score sampling [1, 39, 48] is a more elaborate procedure for selecting the pivots in the Nyström approximation. This method first approximates the ridge leverage scores, which are defined as the diagonal entries $\bm{\ell}^{\lambda}=\operatorname{diag}(\bm{A}(\bm{A}+\lambda\mathbf{I})^{-1})$ for a parameter $\lambda>0$. Then, the method randomly generates a Nyström approximation whose pivots are selected with probabilities proportional to the ridge leverage scores. While RLS sampling can lead to an approximation that is nearly as accurate as RPCholesky (see Figure˜1), it can also produce lower-quality approximations given challenging inputs [13]. Additionally, RLS sampling is slower than the blocked implementation of RPCholesky given in Algorithm˜4, because of the requirement to approximate the ridge leverage scores. For example, when we apply RLS sampling with the RecursiveRLS algorithm [39] to compute a rank-$1000$ approximation of a kernel matrix of size $N=4\times 10^{4}$, the runtime is $2.3\times$ slower than RPCholesky’s runtime.
More expensive Gaussian preconditioner.
@@
$\bm{\widehat{A}}=\bm{A}\bm{Z}(\bm{Z}^{*}\bm{A}\bm{Z})^{\dagger}\bm{Z}^{*}\bm{A},$
where $\bm{Z}\in\real^{N\times r}$ is a matrix with independent standard normal entries. Forming the Gaussian Nyström approximation requires $r$ matrix–vector multiplications with the full kernel matrix. To maintain a cost of $\mathcal{O}(N^{2})$ operations, we can only run Gaussian Nyström with a constant approximation rank $r=\mathcal{O}(1)$, whereas we can run RPCholesky preconditioning with a much larger approximation rank $r=\mathcal{O}(\sqrt{N})$. Because of the higher approximation rank, RPCholesky preconditioning leads to a more accurate approximation with stronger guarantees.
Other approaches.
We briefly mention two other types of approximations that can be used to build preconditioners for the full-data KRR problem. First, certain types of kernel matrices can be approximated using random features [17, 5], but the quality of the approximation tends to be poor. Numerical tests indicate that the random features approach does not yield a competitive preconditioner [17, 59]. See Figures˜1 and 2 for further numerical comparisons.
Second, kernel matrices can be approximated using carefully constructed hierarchical approximations [17, 3, 12, 60] or sparse inverse approximations [50, 61, 28]. These approaches extend beyond low-rank structure and provide richer matrix approximation classes; however, the applicability to high-dimensional data sets remains unclear [3]. So far, these approaches have only been theoretically validated for low-dimensional input data.